Hastings Corporation is interested in acquiring Vandell Corporation. Vandell has 1 million shares outstanding and a target capital structure consisting of 30% debt; its beta is 1.50 (given its target capital structure). Vandell has $9.48 million in debt that trades at par and pays an 7.7% interest rate. Vandell’s free cash flow (FCF0) is $1 million per year and is expected to grow at a constant rate of 6% a year. Both Vandell and Hastings pay a 30% combined federal and state tax rate. The risk-free rate of interest is 3% and the market risk premium is 6%. Hastings Corporation estimates that if it acquires Vandell Corporation, synergies will cause Vandell’s free cash flows to be $2.5 million, $3.1 million, $3.4 million, and $3.53 million at Years 1 through 4, respectively, after which the free cash flows will grow at a constant 6% rate. Hastings plans to assume Vandell’s $9.48 million in debt (which has an 7.7% interest rate) and raise additional debt financing at the time of the acquisition. Hastings estimates that interest payments will be $1.6 million each year for Years 1, 2, and 3. After Year 3, a target capital structure of 30% debt will be maintained. Interest at Year 4 will be $1.442 million, after which the interest and the tax shield will grow at 6%. Indicate the range of possible prices that Hastings could bid for each share of Vandell common stock in an acquisition. Round your answers to the nearest cent. Do not round intermediate calculations. The bid for each share should range between $ per share and $ per share.
Answer #1
Calculation of WACC to discount Vandell's future cash flows:
Cost of Equity
as per CAPM
Ke=RFR+(Beta*Mkt. Risk premium)
ie. Ke=3%+(1.5*6%)=
12%
After-tax cost of debt=
7.7%*(1-30%)=
5.39%
So, WACC=(We*Ke)+(Wd*Kd)
ie.(70%*12%)+(30%*5.39%)=
10.02%
Before synergy
Value of Vandell using
Constant-growth FCF model
by discounting at the above WACC (10.02%) & finding its PV of future cash flows
ie.
FCF1/(r-g)
(FCF0*(1+g))/(r-g)
Where,
g= the constant growth rate ,r= 6%
& r= the reqd. return or WACC=10.02%
Applying the values,
1000000*1.06/(10.02%-6%)=
26368159
So,Firm Value of Vandell before considering synergy= $ 26368159
Value of Debt = 9480000
So, value of Equity=
26368159-9480000=
16888159
No.of equity shares o/s = 1000000
So, Value /share=
16888159/1000000=
16.89
Subsequent to synergy:
Equity Value of Vandell can be straight-away found out
by discounting the following to their Present Values
FCFs of Yrs. 1,2,3,4 & PV of terminal cash flow
Less: PV of all additional debt- interest payments (incl.terminal value)
Add:PV of additional debt-interest Interest tax shields that will be available because of synergies
Year 0 1 2 3 4
FCF 2.5 3.1 3.4 3.53
Terminal FCF(3.53*1.06)/(10.02%-6%) 93.0796
Less: Interest payments -1.6 -1.6 -1.6 -1.442
Less: Terminal Int.exp.(1.442*1.06)/(10.02%-6%) -38.0229
Add: Interest tax shields(Int.exp.*Tax rate) 0.48 0.48 0.48 0.4326
Add: Terminal Int.Tax Shield 11.40687
FCF to Equity/FCFE 1.38 1.98 2.28 68.98418
PV F at 10.02%(1/1.1002^n) 0.908926 0.826146 0.750905 0.682517
PV at 10.02% 1.254317 1.635769 1.712064 47.08287
NPV or Value to Equity 51.685023
ie.Value of Vandell's Equity after synergy= 51685023
No.of equity share o/s 1000000
Value/share 51.69
Value before synergy 16.89
Value after synergy 51.69
So, the range of possible prices that Hastings could bid for each share of Vandell common stock in an acquisition=
$ 16.89 to $ 51.69